# Morse-Bott inequalities, Topology Change and Cobordisms to Nothing

### v1

@@ 4 Some more physical examples
After the theoretical introduction and mathematical results from the previous sections, we will now consider some relatively realistic constructions, as well as the qualitative properties of the cobordisms to nothing/end of the world branes.
@@ 4.1 Spheres with fluxes
We first begin by considering a pretty straightforward examples. Take for this the 10d Einstein-dilaton-RR sector of the type II bosonic supergravity action in Einstein frame Ibanez:2012zz ; VanRiet:2023pnx ,
$S_{\rm II}\supseteq\int\mathrm{d}^{10}x\sqrt{-G_{10}}\left\{\mathcal{R}_{10}-(\partial\hat{\Phi})^{2}-\frac{1}{2n!}e^{\frac{5-n}{\sqrt{2}}\hat{\Phi}}|F_{n}|^{2}\right\}\;,$ (4.1)
where $\hat{\Phi}$ is the canonically normalized 10 dimensional dilaton, and for simplicity we have considered a single RR field strength. One can compactify the above theory on a $n$-sphere $\mathbb{S}^{n}$ with volume $\mathcal{V}$ in $M_{\rm Pl,10}$ units and $N$ units of flux, which assuming that $F_{p}$ only has internal legs, results in a scalar potential in the $(d=10-n)$-dimensional EFT from the curvature and flux contributions:
@@
where $\rho=\sqrt{\frac{8}{(10-d)(d-2)}}\log\frac{\mathcal{V}}{\mathcal{V}_{0}}$ is the canonically normalized radion, with $\mathcal{V}_{0}=\left(\frac{\kappa_{10}}{\kappa_{d}}\right)^{2}$ some fixed volume. While for $d=5$ (precisely when $\Phi$ is not sourced by the RR field strength, see (4.1)) (4.2) only depends on the volume $\mathcal{V}\sim N^{\frac{5}{4}}$ and has an AdS minimum $V_{\rm AdS}\sim-N^{-\frac{4}{3}}$ which becomes deeper as $N$ decreases, for $d\neq 5$ the potential gradient flow pushes towards large volume configurations and small/large coupling (depending on the sign of $5-d$), with $\hat{\Phi}=\sqrt{\frac{d-2}{10-d}}\frac{d-9}{5-d}\rho$.
Now, as explained in section 3.3, the inclusion of $(p-1)$-forms in our theory (which act as gauge potentials for the RR field strength) results in a conserved topological charge (the flux number) upon compactification on oriented surfaces such as $\mathbb{S}^{p}$. In order to kill the $\Omega_{p}^{SO,\,A_{p-1}}$ invariant, D$(8-p)$-brane defects must be included. Precisely this D$(8-p)$-brane will act as the EotW brane, with the $\mathbb{S}^{p}$ sphere shrinking as we approach it. This was studied in more generality in Angius:2022aeq , to which we refer for more detail. In general the inclusion of a D-brane results in a singular term in the EFT action, and as such the bordism $\mathcal{B}_{p+1}$ will have a conical singularity at its tip, where the Dp-brane is located. However, we still have $\mathcal{B}_{p+1}\simeq\mathbb{D}^{p+1}$, and so we can still consider a smooth structure on the bordism and the results from section 3.
In the presence of $F_{n}$ RR flux, a D$(8-n)$ brane (with $n=2,\dots,7$) [^12] We do not consider D8- and D7-branes as their backreaction prevents asymptotically flat spacetime, complicating the analysis. solution to (4.1) has the following 10d Einstein metric and $\hat{\Phi}$ profiles Ibanez:2012zz :
@@
Upon compactification on $\mathbb{S}^{n}$ (transversal to the brane) down to $d=10-n$ dimensions, the D$(8-n=d-2)$-brane has now spatial codimension one (as expected from EotW branes), with the Einstein frame metric and the radion profile now given by Angius:2022aeq
@@
with $V(r)\to 0$ as $r\to 0$ except for $d=5$ (equivalently $n=5$ or D3 branes), where the volume remains finite, and indeed the D$(d-2)$-branes serve as EotW branes, coinciding with the tip of the $\mathbb{S}^{10-d}$ bordism to nothing, topologically given by the solid disk $\mathbb{D}^{11-d}$. The $\Omega_{d-2}^{SO,\,A_{10-d}}$ defect given by the D-brane is precisely located in the topology change (in this case into nothing).
For completion, we refer to Angius:2022aeq , where these solutions were studied with the EotW brane formalism, with the following critical exponent [^14] This critical exponent can also be immediately computed by obtaining the exponential rate of potential (4.2) along the trajectory $\hat{\Phi}=\sqrt{\frac{d-2}{10-d}}\frac{d-9}{5-d}\rho$ the scalars follow for $d\neq 5$.
@@
The above description only discusses the local EotW behavior of the scalars as they approach the brane/cobordism wall and the sphere threaded by fluxes shrinks to a point, which as we have argued for and seen can happen without intermediate topology changes. Actual vacuum instability solutions corresponding to BoN are much more involved, and we will not detail them. For this we refer to the literature, such as Yang:2009wz ; Blanco-Pillado:2010xww ; Blanco-Pillado:2010vdp ; Brown:2010mf ; Blanco-Pillado:2016xvf ; Draper:2023ulp ; Blanco-Pillado:2023aom .
@@ 4.2 Type IIB string theory on compact Riemann surfaces
We will now consider a more involved example with actual intermediate topology changes being needed in order to have a smooth(able) bordism to nothing, and study how this intermediate changes in the internal dimension affect the lower-dimensional EFT. The most straightforward construction for this is Type IIB string theory compactified on a Riemann surface $\Sigma_{g}$ of genus $g\geq 2$.
For simplicity, we will not consider the full ${\rm Spin}-GL^{+}(2,\mathbb{Z})$ bordism of type IIB string theory [^15] Type IIB string theory has fermions and a duality group $SL(2,\mathbb{Z})$, which after including the F-theory torus reflections (and their action on fermions) is promoted to $GL^{+}(2,\mathbb{Z})$ Tachikawa:2018njr ; Pantev:2016nze . This way, the actual structure our bordism must take into consideration is ${\rm Spin}-GL^{+}(2,\mathbb{Z})$ (see Debray:2021vob ; Dierigl:2022reg ; Debray:2023yrs for more on this, including the rest of cobordism groups and generators, as well as properties of the required defects), with $\Omega^{{\rm Spin}-GL^{+}(2,\mathbb{Z})}_{1}=\langle\{\mathbb{S}^{1}_{R},L^{1}_{4}\}\rangle\simeq\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}\;,\qquad\Omega^{{\rm Spin}-GL^{+}(2,\mathbb{Z})}_{2}=\langle\mathbb{S}^{1}_{p}\times\mathbb{S}^{1}_{R}\rangle\simeq\mathbb{Z}_{2}\;,$ (4.8) where $\mathbb{S}^{1}_{p}$ is the circle with periodic boundary conditions, $\mathbb{S}^{1}_{R}$ is the circle with a non-trivial $GL^{+}(2,\mathbb{Z})$ reflection bundle, and $L^{1}_{4}\simeq\mathbb{S}^{1}/\mathbb{Z}_{4}$ a Lens space, naturally endowed with a $\mathbb{Z}_{4}$ bundle. and simply restrict to Spin bordism, with Anderson1967:SPIN ; McNamara:2019rup
@@
For our purposes, a Riemann surface $\Sigma_{g}$ is determined by its genus $g$ and its Spin structure. Having no torsion, homology is set by the Betti numbers $\vec{b}(\Sigma_{g})=(1,2g,1)$, so there are $b_{1}(\Sigma_{g})=2g$ 1-cycles, paired in intersecting dual cycles. By using (3.8) and (3.9), the Poincaré polynomial of the bordism $\mathcal{B}_{3}$ is $\mathsf{P}_{t}(\mathcal{B}_{3})=1+gt$ [^16] Riemann surfaces automatically set the Betti numbers of their bordism to nothing (assuming it has a single connected component), as the Half dies, half lives theorem implies $b_{1}(\mathcal{B}_{3})=\frac{1}{2}b_{1}(\Sigma_{g})=g$. Because of the low dimensionality we are working with, the only possible critical submanifolds are $\{p\}$ and $\mathbb{S}^{1}$, of dimension $c=0$ and $c=1$, respectively. Now, as at each critical locus with index $\lambda$ we have a $(\lambda+c-1)$-cycle shrinking to nothing, this only allows us to have (a) a 1-cycle collapsing on a point, (b) a 1-cycle collapsing on $\mathbb{S}^{1}$, or (c) a 2-cycle without internal cycles [^17] If torsion has not been induced on $\mathcal{B}_{3}$, this is then $\mathbb{S}^{2}$ shrinking. collapsing on a point. As the two last possibilities saturate the $\lambda+c-1\leq 2=\dim\Sigma_{g}$ inequality, and thus correspond to the final topology change into nothing, there can only be one of them.
@@
with $\mathsf{R}(t)\in\mathbb{Z}_{\geq 0}[t]$ and $b+c=1$. In order to have $1+t$ being a factor of the LHS of the above equation, it is needed that $a=g+c-1$. This results in the following options, depending on the genus $g$:
@@
  • $g\geq 2$: The contribution of $a=g+c-1$ to the total number of topology changes is minimized by $(a,b,c)=(g-1,1,0)$. As each (a)-type topology change “kills” a pair of dual 1-cycles, this results in a final $\Sigma_{1}=\mathbb{T}^{2}$ which shrinks through a type-(b) topology change, with a total $N=g$. An extra (a) and (c) can be in place of the last (b), amounting for $N=g+1$.
This way, a bordism (that does not increment the number of internal components or 1-cycles) between $\Sigma_{g}$ and nothing features either $g$ ($g-1$ type (a) and a final (b)) or $g+1$ ($g$ type (a) and a final (c)) topology changes. At each of these the compact manifold will change as we move in the $\rho$ direction of the bordism, resulting in different low-energy EFTs behind each domain wall.
In Appendix A some basics about Riemann surfaces and their shape moduli are presented. While the final shrinking of $\mathbb{T}^{2}$ or $\mathbb{S}^{2}$ (type (b) or (c) topology changes from the above description) is well understood, accounting simply for a compactification limit in the EotW brane description, type (a) changes, with a 1-cycle pinching on a point and decaying, is not as simple. This can be well understood through tachyon condensation from a world-sheet perspective Saltman:2004jh ; Adams:2005rb . In Appendix B we give a brief overview of how 1-cycles of Riemann surfaces with anti-periodic boundary conditions decay once their thickness at some point of the handle becomes small enough, with tachyon condensation effectively preventing spacetime degrees of freedom from propagating along the handle direction.
@@
After the needed introductory remarks, we can proceed to study the changes in the lower-dimensional EFT under the topology changes. To do so, we first take the 10D Type IIB SUGRA bosonic effective action in Einstein frame Ibanez:2012zz ; VanRiet:2023pnx
@@
again considering only the bosonic part. We now compactify our theory on a Riemann surface $\Sigma_{g}$, with volume $\mathcal{V}$ in 10d Planck units. Taking the legs of $F_{1}$ and one of $H_{3}$ lay on $\Sigma_{g}$, the following 8d effective action results:
@@
[…]
@@
where $\mathbb{A}_{AB}(\vec{\tau})$ is a positive definite symmetric matrix with $\det\mathbb{A}=1$, given in terms of the period matrix $\Omega$, depending only on the complex structure moduli $\vec{\chi}$. It is given, as derived in Appendix A, by (A.18), Saltman:2004jh ; Adams:2005rb
@@
As for the complex structure moduli, these are stabilized for a generic flux configuration of $F_{1}$ Saltman:2004jh . [^18] To see this, notice that all the eigenvalues $\mathcal{A}_{AB}\propto\mathbb{A}_{AB}$ are positive. From the definition (A.7) of the period matrix $\Omega^{ij}$, in a topology change $\Sigma_{g}\to\Sigma_{g-1}$ a 1-cycle pinches and becomes small, and as such some eigenvalue of ${\rm Im}\Omega=\frac{1}{2i}(\Omega-\bar{\Omega})$ goes to 0, or equivalently, one of $(\Omega-\bar{\Omega})^{-1}$ blows up. Then in such a limit $\mathbb{A}(\tau)$ has some (positive) eigenvalue diverging. An arbitrary flux configuration is expected to have $F_{1}$ components along the complete $\mathbb{A}(\tau)$ eigenbasis, so that for any diverging eigenvalue the flux part of $V(\varphi)$ diverges. As this happens for any 1-cycle pinching and the flux potential is non-negative, there must be some minimum in the bulk of the complex structure moduli space, resulting in their stabilization. Note that, as explained in Appendix B, for the topology change to occur it is not needed that the cycle radius actually pinches to zero, as for $\mathsf{R}\leq\sqrt{2\alpha^{\prime}}$ (with $\alpha^{\prime}$ the string tension) a tachyonic instability appears. This means that the $\Sigma_{g}\to\Sigma_{g-1}$ transition does not encounter an infinite potential barrier, but rather a finite-height slope up-hill until the instability region is reached and tachyon condensation occurs.
@@
Figure 6: Sketch of different topology changes along the cobordism to nothing, depicting the intermediate losing of a handle for $g\geq 2$ (including the arrangement of the D7-domain walls, in yellow, and the D$(-1)$-$\overline{\text{D}(-1)}$ bounded on it, in red with a dashed line presenting the 1-chain joining them) and the two possible final topology changes into nothing, $\mathbb{T}^{2}\to\emptyset$ and $\mathbb{S}^{2}\to\emptyset$. The qualitative behavior of the scalar potential is also depicted.
@@
By inspection of action (4.14), it is evident that compactifying on $\Sigma_{g}$ results in distinct low energy EFTs depending on the value of $g$, with the topology changes loci acting as domain walls between different theories. We will now comment on these changes occur as we move in the radial direction towards the end of the bordism to nothing. Assuming that the different topology changes are sufficiently separated from one another such that a lower $d=8$ description is valid between them,
@@
  • The most evident feature of $\Sigma_{g}\to\Sigma_{g-1}$ is the change in the scalar potential $V(\vec{\varphi})$, as depicted in Figure 6. We first take the internal curvature term. Gauss-Bonnet theorem applied to compact Riemann surfaces result in this being proportional to $(g-1)$ times an extra $\mathcal{V}^{-4/3}$ factor coming from the rescaling to 8d Einstein frame. For $g\geq 2$ $\Sigma_{g}$ are negatively curved, resulting in a runaway positive potential to large volume. At the topology change locus a cycle pinches, with the overall volume of the internal manifold not being dramatically affected. Assuming that said volume remains approximately constant as we cross the domain wall (if our bordism $\mathcal{B}_{3}$ is smooth then this is the case), then the internal curvature contribution to the vacuum energy experiences a change of $\Delta V_{\mathcal{R}}\approx-\pi\left(\mathcal{V}/\mathcal{V}_{0}\right)^{-4/3}<0$, with the change being energetically favored. For $g=1$ (i.e. a torus compactification), there is no internal contribution to the curvature. If our bordism features an additional topology change to the sphere $\Sigma_{0}=\mathbb{S}^{2}$, the curvature potential becomes negative. In this case the $F_{1}$ flux potential is 0, as there are no more 1-cycles, so there is no extra contributions to the potential (4.2). This drives $\mathcal{V}\to 0$, corresponding precisely to the $\mathbb{S}^{2}\to\emptyset$ final topology change into nothing.
@@
    As for the flux contributions to the potential, apart from the overall $\mathcal{V}^{-4/3}e^{\Phi}$ factor driving towards a weak coupling and large volume limit there is the $\sum_{I=1}^{2g}Q^{I\intercal}\mathbb{A}(\vec{\chi})Q^{I}$ factor. We can take the factorization approximation so that from (4.18),
@@
    As for the units of flux lost in the topology change process, we must differentiate between those associated with the shrinking cycle $c$ ($Q$) and those with its dual $c$ ($Q^{\prime}$). As was explained in sections 3.2 and 3.3, the shrinking of the cycle can be understood as a cobordism into nothing, with $\mathbb{S}^{1}=\partial\mathbb{D}^{2}$. In order for this to be possible, we need a defect at the tip of the disk which eats up the charge. As reasoned in section 3.3, this corresponds to a D7-brane, magnetically charged under the $C_{8}$ form dual to $C_{0}$. The (6+1)-dimensional topology change locus/domain wall (precisely where the flux potential naively would blow up) corresponds to the D7 brane, with the extra direction wrapped along the dual 1-cycle. This fixes the position of this intermediate D7 to be precisely at the topology change locus.
@@
    We argue that this corresponds to a pair of $\rm D(-1)-\overline{D(-1)}$ branes (D-instantons) localized on the D7 Billo:2021xzh ; Reymond:2024mwe , that are unstable and eventually annhilate each other, along with their charges. As a result these charges are not present at the other side of the domain wall, and do not contribute to the potential anymore.

### v2

@@ 4 Some more physical examples
After the theoretical introduction and mathematical results from the previous sections, we will now consider some more realistic constructions, as well as the qualitative properties of the cobordisms to nothing/end of the world branes.
@@ 4.1 Spheres with fluxes
We first begin by considering a pretty straightforward set of examples. Take for this the 10d Einstein-dilaton-RR sector of the type II bosonic supergravity action in Einstein frame Ibáñez and Uranga (2012); Van Riet and Zoccarato (2024),
$S_{\rm II}\supseteq\frac{1}{2\kappa_{10}^{2}}\int\mathrm{d}^{10}x\sqrt{-G_{10}}\left\{\mathcal{R}_{10}-(\partial\hat{\Phi})^{2}-\frac{1}{2n!}e^{\frac{5-n}{\sqrt{2}}\hat{\Phi}}|F_{n}|^{2}\right\}\;,$ (4.1)
where $\hat{\Phi}$ is the canonically normalized 10 dimensional dilaton, and for simplicity we have considered a single RR field strength. One can compactify the above theory on a $n$-sphere $\mathbb{S}^{n}$ with volume $\mathcal{V}$ in $M_{\rm Pl,10}$ units and $N$ units of flux, which, assuming that $F_{p}$ only has internal legs, results in a scalar potential in the (d=10n)(d=10-n)-dimensional EFT from the curvature and flux contributions:
@@
where $\rho=\sqrt{\frac{8}{(10-d)(d-2)}}\log\frac{\mathcal{V}}{\mathcal{V}_{0}}$ is the canonically normalized radion, with $\mathcal{V}_{0}=\left(\frac{\kappa_{10}}{\kappa_{d}}\right)^{2}$ some fixed volume. While for $d=5$ (precisely when $\Phi$ is not sourced by the RR field strength, see (4.1)) (4.2) only depends on the volume $\mathcal{V}\sim N^{\frac{5}{4}}$ and has an AdS minimum $V_{\rm AdS}\sim-N^{-\frac{4}{3}}$ which becomes deeper as $N$ decreases, for $d\neq 5$ the potential gradient flow pushes towards small volume configurations and large/small coupling (depending on the sign of 5d5-d), with $\hat{\Phi}=\sqrt{\frac{d-2}{10-d}}\frac{d-9}{5-d}\rho$.
Now, as explained in section 3.3, the inclusion of (p1)(p-1)-forms in our theory (which act as gauge potentials for the RR field strength) results in a conserved topological charge (the flux number) upon compactification on oriented surfaces such as $\mathbb{S}^{p}$. In order to kill the $\Omega_{p}^{SO,\,A_{p-1}}$ invariant, D(8p)(8-p)-brane defects must be included. Precisely this D(8p)(8-p)-brane will act as the EotW brane, with the $\mathbb{S}^{p}$ sphere shrinking as we approach it. This was studied in more generality in Angius et al. (2022a), to which we refer for more detail. In general the inclusion of a D-brane results in a singular term in the EFT action, and as such the bordism $\mathcal{B}_{p+1}$ will have a conical singularity at its tip, where the Dp-brane is located. However, we still have $\mathcal{B}_{p+1}\simeq\mathbb{D}^{p+1}$, and so we can still consider a generic smooth structure on the bordism and the results from section 3.
In the presence of $F_{n}$ RR flux, a D(8n)(8-n) brane (with $n=2,\dots,7$) [^14] We do not consider D8- and D7-branes as their backreaction prevents asymptotically flat spacetime, complicating the analysis. solution to (4.1) has the following 10d Einstein metric and $\hat{\Phi}$ profiles Ibáñez and Uranga (2012):
@@
Upon compactification on $\mathbb{S}^{n}$ (transversal to the brane) down to d=10nd=10-n dimensions, the D(8n=d2)(8-n=d-2)-brane has now spatial codimension one (as expected from EotW branes), with the Einstein frame metric and the radion profile now given by Angius et al. (2022a)
@@
with $V(r)\to 0$ as $r\to 0$ except for $d=5$ (equivalently $n=5$ or D3 branes), where the volume remains finite, and indeed the D(d2)(d-2)-branes serve as EotW branes, coinciding with the tip of the $\mathbb{S}^{10-d}$ bordism to nothing, topologically given by the solid disk $\mathbb{D}^{11-d}$. The $\Omega_{d-2}^{SO,\,A_{10-d}}$ defect given by the D-brane is precisely located in the topology change (in this case into nothing).
For completion, we refer to Angius et al. (2022a), where these solutions were studied with the EotW brane formalism, where the following critical exponent [^16] This critical exponent can also be immediately computed by obtaining the exponential rate of potential (4.2) along the trajectory $\hat{\Phi}=\sqrt{\frac{d-2}{10-d}}\frac{d-9}{5-d}\rho$ the scalars follow for $d\neq 5$.
@@
The above description only discusses the local EotW behavior of the scalars as they approach the brane/cobordism wall and the sphere threaded by fluxes shrinks to a point, which as we have argued for and seen, can happen without intermediate topology changes. Actual vacuum instability solutions corresponding to BoN are much more involved, and we will not detail them. For this we refer to the literature, such as Yang (2010); Blanco-Pillado and Shlaer (2010); Blanco-Pillado et al. (2010); Brown and Dahlen (2011); Blanco-Pillado et al. (2016); Draper et al. (2023); Blanco-Pillado et al. (2024a).
@@ 4.2 Type IIB string theory on compact Riemann surfaces
We will now consider a more involved example with actual intermediate topology changes being needed in order to have a smooth(able) bordism to nothing, and study how this intermediate changes in the internal dimension affect the lower-dimensional EFT. For this we will consider Type IIB string theory compactified on a Riemann surface $\Sigma_{g}$ of genus $g\geq 2$.
For simplicity, we will not consider the full ${\rm Spin}-GL^{+}(2,\mathbb{Z})$ bordism of type IIB string theory [^17] Type IIB string theory has fermions and a duality group $SL(2,\mathbb{Z})$, which after including the F-theory torus reflections (and their action on fermions) is promoted to $GL^{+}(2,\mathbb{Z})$ Tachikawa and Yonekura (2019); Pantev and Sharpe (2016). This way, the actual structure our bordism must take into consideration is ${\rm Spin}-GL^{+}(2,\mathbb{Z})$ (see Debray et al. (2021); Dierigl et al. (2023); Debray et al. (2023) for more on this, including the rest of cobordism groups and generators, as well as properties of the required defects), with $\Omega^{{\rm Spin}-GL^{+}(2,\mathbb{Z})}_{1}=\langle\{\mathbb{S}^{1}_{R},L^{1}_{4}\}\rangle\simeq\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}\;,\qquad\Omega^{{\rm Spin}-GL^{+}(2,\mathbb{Z})}_{2}=\langle\mathbb{S}^{1}_{p}\times\mathbb{S}^{1}_{R}\rangle\simeq\mathbb{Z}_{2}\;,$ (4.8) where $\mathbb{S}^{1}_{p}$ is the circle with periodic boundary conditions, $\mathbb{S}^{1}_{R}$ is the circle with a non-trivial $GL^{+}(2,\mathbb{Z})$ reflection bundle, and $L^{1}_{4}\simeq\mathbb{S}^{1}/\mathbb{Z}_{4}$ a Lens space, naturally endowed with a $\mathbb{Z}_{4}$ bundle. and simply restrict to Spin bordism, with Anderson et al. (1967); McNamara and Vafa (2019)
@@
For our purposes, a Riemann surface $\Sigma_{g}$ is determined by its genus $g$ and its Spin structure. Having no torsion, homology is set by the Betti numbers $\vec{b}(\Sigma_{g})=(1,2g,1)$, so there are $b_{1}(\Sigma_{g})=2g$ 1-cycles, paired in intersecting dual cycles. By using (3.8) and (3.9), the Poincaré polynomial of the bordism $\mathcal{B}_{3}$ is $\mathsf{P}_{t}(\mathcal{B}_{3})=1+gt$. [^18] Riemann surfaces automatically set the Betti numbers of their bordism to nothing (assuming it has a single connected component), as the Half dies, half lives theorem implies $b_{1}(\mathcal{B}_{3})=\frac{1}{2}b_{1}(\Sigma_{g})=g$. Because of the low dimensionality we are working with, the only possible critical submanifolds are $\{p\}$ and $\mathbb{S}^{1}$, of dimension $c=0$ and $c=1$, respectively. Now, as at each critical locus with index $\lambda$ we have a $(\lambda+c-1)$-cycle shrinking to nothing, this only allows us to have (a) a 1-cycle collapsing on a point, (b) a 1-cycle collapsing on $\mathbb{S}^{1}$, or (c) a 2-cycle without internal cycles [^19] If torsion has not been induced on $\mathcal{B}_{3}$, this is then $\mathbb{S}^{2}$ shrinking. collapsing on a point. As the two last possibilities saturate the $\lambda+c-1\leq 2=\dim\Sigma_{g}$ inequality, and thus correspond to the final topology change into nothing, there can only be one of them.
@@
with $\mathsf{R}(t)\in\mathbb{Z}_{\geq 0}[t]$ and $b+c=1$. In order to have $1+t$ being a factor of the LHS of the above equation, it is needed that a=g+c1a=g+c-1. This results in the following options, depending on the genus $g$:
@@
  • $g\geq 2$: The contribution of a=g+c1a=g+c-1 to the total number of topology changes is minimized by (a,b,c)=(g1,1,0)(a,b,c)=(g-1,1,0). As each (a)-type topology change “kills” a pair of dual 1-cycles, this results in a final $\Sigma_{1}=\mathbb{T}^{2}$ which shrinks through a type-(b) topology change, with a total $N=g$. An extra (a) and (c) can be in place of the last (b), amounting for $N=g+1$.
This way, a bordism (that does not increment the number of internal components or 1-cycles) between $\Sigma_{g}$ and nothing features either $g$ (g1g-1 type (a) and a final (b)) or $g+1$ ($g$ type (a) and a final (c)) topology changes. At each of these, the compact manifold will change as we move in the $\rho$ direction of the bordism, resulting in different low-energy EFTs behind each domain wall.
In Appendix A some basics about Riemann surfaces and their shape moduli are presented. While the final shrinking of $\mathbb{T}^{2}$ or $\mathbb{S}^{2}$ (type (b) or (c) topology changes from the above description) is well understood, (accounting simply for a compactification limit in the EotW brane description), type (a) changes, with a 1-cycle pinching on a point and decaying, are not as simple. This can be well understood through tachyon condensation from a world-sheet perspective Saltman and Silverstein (2006); Adams et al. (2005). In Appendix B we give a brief overview of how 1-cycles of Riemann surfaces with anti-periodic boundary conditions decay once their thickness at some point of the handle becomes small enough, with tachyon condensation effectively preventing spacetime degrees of freedom from propagating along the handle direction.
@@
After the above introductory remarks, we can proceed to study the changes in the lower-dimensional EFT under the topology changes. To do so, we first take the 10D Type IIB SUGRA bosonic effective action in Einstein frame Ibáñez and Uranga (2012); Van Riet and Zoccarato (2024)
@@
again considering only the bosonic part. We now compactify our theory on a Riemann surface $\Sigma_{g}$, with volume $\mathcal{V}$ in 10d Planck units. Taking the legs of $F_{1}$ and one of $H_{3}$ laying on $\Sigma_{g}$, the following 8d effective action results:
@@
where here $\vec{\varphi}=(\Phi,\mathcal{V},\vec{\tau})$ are the moduli corresponding to the 10d dilaton, $\Sigma_{g}$ volume and complex structure (respectively 0, 1 and 3(g1)3(g-1) for g=0, 1g=0,\,1 or more), and
$H_{2}^{A}\equiv\mathrm{d}B_{1}^{A}=\int_{a_{A}}H_{3}\;,\qquad A=1,\dots,\,2g$ (4.15)
@@
where $\mathbb{A}_{AB}(\vec{\tau})$ is a positive definite symmetric matrix with $\det\mathbb{A}=1$, given in terms of the period matrix $\Omega$, depending only on the complex structure moduli $\vec{\chi}$. It is given, as derived in Appendix A, by (A.18), Saltman and Silverstein (2006); Adams et al. (2005)
@@
As for the complex structure moduli, these are stabilized for a generic flux configuration of $F_{1}$ Saltman and Silverstein (2006). [^20] To see this, notice that all the eigenvalues $\mathcal{A}_{AB}\propto\mathbb{A}_{AB}$ are positive. From the definition (A.7) of the period matrix $\Omega^{ij}$, in a topology change $\Sigma_{g}\to\Sigma_{g-1}$ a 1-cycle pinches and becomes small, and as such some eigenvalue of ${\rm Im}\Omega=\frac{1}{2i}(\Omega-\bar{\Omega})$ goes to 0, or equivalently, one of $(\Omega-\bar{\Omega})^{-1}$ blows up. Then in such a limit $\mathbb{A}(\tau)$ has some (positive) eigenvalue diverging. An arbitrary flux configuration is expected to have $F_{1}$ components along the complete $\mathbb{A}(\tau)$ eigenbasis, so that for any diverging eigenvalue the flux part of $V(\varphi)$ diverges. As this happens for any 1-cycle pinching and the flux potential is non-negative, there must be some minimum in the bulk of the complex structure moduli space, resulting in their stabilization. Note that, as explained in Appendix B, for the topology change to occur it is not needed that the cycle radius actually pinches to zero, as for $\mathsf{R}\leq\sqrt{2\alpha^{\prime}}$ (with $\alpha^{\prime}$ the string tension) a tachyonic instability appears. This means that the $\Sigma_{g}\to\Sigma_{g-1}$ transition does not encounter an infinite potential barrier, but rather a finite-height slope up-hill until the instability region is reached and tachyon condensation occurs.
@@
Figure 6: Sketch of different topology changes along the cobordism to nothing, depicting the intermediate losing of a handle for $g\geq 2$ (including the arrangement of the D7-domain walls, in yellow, and the D(1)(-1)-$\overline{\text{D}(-1)}$ bounded on it, in red with a dashed line presenting the 1-chain joining them) and the two possible final topology changes into nothing, $\mathbb{T}^{2}\to\emptyset$ and $\mathbb{S}^{2}\to\emptyset$. The qualitative behavior of the scalar potential is also depicted.
@@
By inspection of action (4.14), it is evident that compactifying on $\Sigma_{g}$ results in distinct low energy EFTs depending on the value of $g$, with the topology changes loci acting as domain walls between different theories. We will now comment on how these changes occur as we move in the radial direction towards the end of the bordism to nothing. Assuming that the different topology changes are sufficiently separated from one another such that a lower $d=8$ description is valid between them,
@@
  • The most evident feature of $\Sigma_{g}\to\Sigma_{g-1}$ is the change in the scalar potential $V(\vec{\varphi})$, as depicted in Figure 6. We first take the internal curvature term. Gauss-Bonnet theorem applied to compact Riemann surfaces result in this being proportional to (g1)(g-1) times an extra $\mathcal{V}^{-4/3}$ factor coming from the rescaling to 8d Einstein frame. For $g\geq 2$ $\Sigma_{g}$ are negatively curved, resulting in a runaway positive potential to large volume. At the topology change locus a cycle pinches, with the overall volume of the internal manifold not being dramatically affected. Assuming that said volume remains approximately constant as we cross the domain wall (if our bordism $\mathcal{B}_{3}$ is smooth then this is the case), then the internal curvature contribution to the vacuum energy experiences a change of $\Delta V_{\mathcal{R}}\approx-\pi\left(\mathcal{V}/\mathcal{V}_{0}\right)^{-4/3}<0$, with the change being energetically favored. For $g=1$ (i.e. a torus compactification), there is no internal contribution to the curvature. If our bordism features an additional topology change to the sphere $\Sigma_{0}=\mathbb{S}^{2}$, the curvature potential becomes negative. In this case the $F_{1}$ flux potential is 0, as there are no more 1-cycles, so there is no extra contributions to the potential (4.2). This drives $\mathcal{V}\to 0$, corresponding precisely to the $\mathbb{S}^{2}\to\emptyset$ final topology change into nothing.
@@
    As for the flux contributions to the potential, apart from the overall $\mathcal{V}^{-4/3}e^{\Phi}$ factor driving towards a weak coupling and large volume limit, there is the $\sum_{I=1}^{2g}Q^{I\intercal}\mathbb{A}(\vec{\chi})Q^{I}$ factor. We can take the factorization approximation so that from (4.18),
@@
    As for the units of flux lost in the topology change process, we must differentiate between those associated with the shrinking cycle $c$ ($Q$) and those with its dual $c$ ($Q^{\prime}$). As was explained in sections 3.2 and 3.3, the shrinking of the cycle can be understood as a cobordism into nothing, with $\mathbb{S}^{1}=\partial\mathbb{D}^{2}$. In order for this to be possible, we need a defect at the tip of the disk which eats up the charge. As reasoned in section 3.3, this corresponds to a D7-brane, electrically charged under the $C_{8}$ form dual to $C_{0}$. The (6+1)-dimensional topology change locus/domain wall (precisely where the flux potential naively would blow up) corresponds to the D7 brane, with the extra direction wrapped along the dual 1-cycle. This fixes the position of this intermediate D7 to be precisely at the topology change locus.
@@
    We argue that this corresponds to a pair of $\rm D(-1)-\overline{D(-1)}$ branes (D-instantons) localized on the D7 Billò et al. (2021); Reymond et al. (2024), that are unstable and annihilate each other, along with their charges. As a result these charges are not present at the other side of the domain wall, and do not contribute to the potential anymore.